# DPImageBench: A Unified Benchmark for Differentially Private Image Synthesis

### v2

@@ Abstract
To address the issue, this paper introduces DPImageBench for DP image synthesis, with thoughtful design across several dimensions: (1) Methods. We study twelve prominent methods and systematically characterize each based on model architecture, pretraining strategy, and privacy mechanism. (2) Evaluation. We include nine datasets and seven fidelity and utility metrics to thoroughly assess these methods. Notably, we find that the common practice of selecting downstream classifiers based on the highest accuracy on the sensitive test set not only violates DP but also overestimates the utility scores. DPImageBench corrects for these mistakes. (3) Platform. Despite the wide variety of methods and evaluation protocols, DPImageBench provides a standardized interface that accommodates current and future implementations within a unified framework. With DPImageBench, we have several noteworthy findings. For example, contrary to the common wisdom that pretraining on public image datasets is usually beneficial, we find that the distributional similarity between pretraining and sensitive images significantly impacts the performance of the synthetic images and does not always yield improvements. In addition, adding noise to low-dimensional features, such as the high-level characteristics of sensitive images, is less affected by the privacy budget compared to adding noise to high-dimensional features, like weight gradients. The former methods perform better than the latter under a low privacy budget, like . The source code is available.111https://github.com/2019ChenGong/DPImageBench.
1 Introduction
Privacy-preserving synthetic data aims to generate artificial data that maintains the properties of real data, enabling data sharing within and across organizations while minimizing privacy risks [1, 2]. Differentially private (DP) dataset synthesis [1, 3] provides a theoretical guarantee to quantify privacy leakage of real data using the synthetic dataset. In recent years, there has been significant development of DP dataset synthesis methods across various domains, including images [4, 5, 6], tabular data [7, 8], text [9, 10, 11], networking data [12, 13, 14] and more.
Among them, DP image synthesis has witnessed impressive progress [5, 6, 15, 16, 4, 17, 18, 19] but lacks a comprehensive understanding of the strengths and weaknesses of existing synthesizers. While there are several studies summarizing advances in this field, they either focus on discussions [20, 1] and do not perform much evaluation, or only cover subfields [21, 22]. This paper focuses on benchmarking DP image dataset synthesis, addressing several challenges:
@@ 1 Introduction
Evaluation: The lack of unified metrics makes it difficult to evaluate the quality of synthetic images. Existing methods such as PE [5] and PrivImage [4] both utilize FID [23] for fidelity evaluation. However, FID does not assess whether the generated images align with human perception, as described in Section 5.3.2. Even worse, the standard practice that evaluates synthetic image quality by tuning classifiers on the sensitive test set to select optimal parameters (used in several methods [19, 15]) is incorrect: they both violate DP and inflate the results.
@@
Contributions. Without a clear understanding of existing methods and an easy-to-use benchmarking platform, it is challenging for the community to make meaningful and reproducible advancements in this field. To bridge this gap, we establish DPImageBench, a benchmark specifically focused on DP image synthesis. DPImageBench addresses the above challenges by making the following contributions:
@@
Taxonomy. We taxonomize methods from two orthogonal dimensions, covering twelve state-of-the-art DP image synthesis algorithms. (1) Private Data Perspective. Based on where noise is added to satisfy DP, we categorize methods into (a) Input-level: adding noise to (some format of) the sensitive images; (b) Model-level: incorporating noise into the gradient or loss during model training; and (c) Output-level: introducing noise during the synthesis process. (2) Public Data Perspective. Existing works [4, 18] use publicly available data from open platforms without privacy concerns to enhance DP image synthesis. We categorize them into three types based on how the public data is utilized: leveraging pretrained models, public datasets without selection, and public datasets with selection. Besides, we argue that the state-of-the-art method DP-Promise [24] does not strictly satisfy DP in Appendix E to prevent others from wrongly using it.
@@
Platform. We establish a platform, DPImageBench, with the following contributions: (1) Fair comparison. DPImageBench standardizes the public images, model architectures, and hyper-parameters to ensure fair comparisons. DPImageBench offers an interface to pretrain the synthesizer, even for the ones that were not originally designed to use a public dataset for pretraining. (2) Straightforward implementation. DPImageBench features an abstract interface compatible with any synthesis algorithm. With its modular design, DPImageBench provides a flexible tool that enables users to select public and sensitive datasets, and to integrate new algorithms by adding functional code to the relevant modules. Since DPImageBench isolates the parameter configurations from implementations, users typically only need to modify the configuration files to tune algorithms. (3) Open Source. DPImageBench is open-sourced. It serves as a starting point for the implementation and evaluation of future methods as well.
@@
Evaluation. We have conducted comprehensive evaluations, and some key highlights include: (1) Datasets. DPImageBench implements synthesis methods on seven sensitive datasets and two public dataset. (2) Fidelity. We establish metrics to assess synthetic images based on similarity, diversity, novelty, and alignment with human perception [25]. (3) Utility. We correct the common evaluation mistakes of using sensitive test images for hyper-parameter tuning. Instead, DPImageBench adopts the standard of using distinct training, validation, and test sets for utility evaluations. Downstream classifier training is performed on the training set, hyper-parameter selection is done on the validation set, and final results are reported on the test set. DPImageBench provides two choices of validation set: (1) the synthetic images, and (2) images having the same distribution as the test set but not overlapping with it. In the latter, DPImageBench adds noise to the validation results to satisfy DP, which is detailed in Section 5.3.4.
@@
Insights to new techniques. The proposed taxonomies and straightforward implementations enable us to analyze the potential of combining different methods and bring insight into developing new technology. We consider some methods mostly in isolation, though they can be effectively combined. For example, the public dataset selection method from PrivImage [4] and creating shortcuts for DP training in DP-FETA [26], as described in Section 4, can be paired with other methods. We present details in Section 7.4.
@@
Pretraining with public datasets does not always benefit DP image synthesis. The effectiveness of pretraining depends heavily on the choice of public dataset and the pertaining strategy (i.e., conditional or unconditional pretraining). Conditional pretraining yields better synthetic performance than unconditional pretraining. A pretraining dataset similar to the sensitive images can enhance synthetic performance.
The performance of adding noise to low-dimensional features, such as high-level characteristics of sensitive images in DP-MERF [28] and training loss in DP-Kernel [17], is less sensitive to the privacy budget compared to adding noise to high-dimensional features, like weight gradients using DP-SGD. The former methods perform better than the latter under a low privacy budget.
The combinations of algorithmic improvements, such as integrating DP-FETA [26] into other methods, do not consistently enhance synthetic performance. This inconsistency underscores the need for further research to effectively leverage the strengths of diverse algorithms.
2 Preliminaries
2.1 Privacy Concepts and Mechanisms
Algorithm | Privacy notion |
DP-MERF [28] | Replace-one |
DP-NTK [29] | Replace-one |
DP-Kernel [17] | Replace-one |
PE [5] | Add-or-remove-one |
GS-WGAN [15] | Both |
DP-GAN [30] | Add-or-remove-one |
DPDM [31] | Add-or-remove-one |
DP-FETA [26] | Add-or-remove-one |
PDP-Diffusion [32] | Add-or-remove-one |
DP-LDM [18] | Add-or-remove-one |
DP-LoRA [19] | Add-or-remove-one |
PrivImage [4] | Add-or-remove-one |
DPImageBench concentrates on image synthesis under differential privacy (DP) constraint. DP [33] provides mathematical guarantees that quantify and limit the amount of information one can derive from the algorithm’s output and has several properties, including composability [34].
Definition 1 (()-DP [33]).
@@ 2.1 Privacy Concepts and Mechanisms
The privacy budget quantifies how much information can be exposed by the algorithm . A smaller value of implies that the algorithm provides stronger privacy guarantees. The can be intuitively understood as failure probability [33]. Given two datasets and , they are considered neighbors (denoted as ) if we can obtain one from the other by adding, removing, or replacing a single image. We define the sensitivity (intuitively, a stability measure) of a function as . As presented in Table I, two widely used definitions for neighboring datasets exist in existing DP image synthesis: the ‘replace-one’ (bounded DP) and ‘add-or-remove-one’ (unbounded DP) definitions. Specifically, given two datasets and , in the ‘replace-one’ definition, neighboring datasets are defined by replacing one image sample, i.e., for image samples and . In the ‘add-or-remove-one’ definition, neighboring datasets are constructed by adding or removing a single image sample, e.g., . We show in Appendix A that all methods that used the replace-one definition can be adapted to add-or-remove-one without affecting their privacy guarantees. As a result, all methods can be compared fairly under the same privacy notion.
@@
represents a multi-dimensional random variable drawn from an i.i.d. normal distribution with a mean of 0 and a variance of , where is a hyper-parameter known as the noise multiplier [33].
DP-SGD. To train machine learning models with DP guarantees, DP-SGD [35] extends standard SGD by controlling the gradient contributions from individual data and applying the Gaussian mechanism. Specifically, in each training iteration, DP-SGD samples a mini-batch from the whole training data and bounds the sensitivity of each gradient using a clipping operation: where indicates the original gradient and is a hyper-parameter. The ‘’ operation scales the gradient norm to be at most , and the model parameters are then updated using the following equation:
where is the learning rate, denotes the loss function of an arbitrary machine learning algorithm, and refers to a mini-batch sampled from the dataset using Poisson sampling with a rate (with indicating its size). The denotes the gradient computed for the sample . Besides, represents i.i.d. Gaussian noise with variance .
As mentioned above, DP-SGD uses only one batch of the sensitive dataset for each iteration of the model parameter update. However, the training process involves many iterations of DP-SGD to ensure that the target model converges. Composition theorems (privacy accounting), e.g., Rényi DP [36], can be used to account for the privacy budget for each iteration and derive the final values of [36].
2.2 Image Generative Models
We introduce three widely used image synthesizers: (1) diffusion models (DMs), including standard DMs [37] and latent DMs [38], (2) Generative Adversarial Networks (GANs) [39], and (3) foundational APIs [5].
Diffusion Models. DMs [37] consist of two processes: (1) The forward diffusion process that progressively adds noise to a clean image until it becomes close enough to pure random noise, generating a sequence of increasingly noisy images , where is the number of noising steps. (2) The reverse process that progressively denoises a random noise to a clean image via a trainable neural network. In the forward diffusion process, the transition between adjacent noisy images, , follows a multi-dimensional Gaussian distribution,
@@ 2.2 Image Generative Models
The final objective of DMs is to train a network to predict the added noise at each step, as defined in [37].
@@
Latent Diffusion Models (LDMs). LDMs [38] are a variant of DMs that operate in a latent space instead of directly in pixel space. LDMs first use a pretrained autoencoder to map images into a lower-dimensional latent representation. The same diffusion and denoising steps discussed before proceed in this compressed latent space rather than on raw pixels. After performing the reverse diffusion steps in latent spaces, the decoder reconstructs final high-resolution images. LDMs operate in a lower-dimensional space, enabling more efficient training and sampling than standard pixel-based DM.
Stable diffusion is a representative example of LDMs that generate images from textual descriptions or class labels. DP-LDM [18] and DP-LoRA [19] leverage stable diffusion [38] as the synthesizer.
Generative Adversarial Nets (GAN). GAN consists of two networks: a generator Gen and a discriminator Dis [39, 40]. The generator Gen receives a random noise vector and outputs an image. The discriminator Dis receives an image and outputs a score, which indicates how real the input image is. The Gen is trained to deceive the discriminator, and the Dis is trained to distinguish whether its input image comes from the true dataset or is generated by the Gen. Both networks are trained simultaneously, constantly improving in response to each other’s updates. Overall, the objective function of GAN is defined as,
@@
Foundational APIs. Foundation APIs are either cloud-based services (e.g., Dall-E [41]) or local software libraries (e.g., Stable Diffusion [38]) that provide simple API access to powerful, pretrained models capable of generating high-quality synthetic images. The underlying model can be any generative model, including diffusion models and GANs.
2.3 DP Image Synthesis
Various sensitive image datasets, such as medical images [42] and face images [43], face the risk of privacy leakage. Attackers may infer information about the training dataset from the synthetic images [44]. DP image synthesis aims to generate artificial images that closely resemble real data while ensuring the privacy protection of the original dataset. DP quantifies leakage to infer the training dataset using the synthetic dataset. These methods enable organizations to share and use synthetic images for various downstream analyses, reducing privacy concerns. We introduce current state-of-the-art DP image synthesis methods and present their taxonomy in Section 4.
3 Overview of DPImageBench
We introduce DPImageBench, a modular toolkit designed to evaluate the performance of DP image synthesis algorithms in terms of utility and fidelity. We envisage DPImageBench for the following four purposes.

### v3

@@ Abstract.
To address the issue, this paper introduces DPImageBench for DP image synthesis, with thoughtful design across several dimensions: (1) Methods. We study twelve prominent methods and systematically characterize each based on model architecture, pretraining strategy, and privacy mechanism. (2) Evaluation. We include nine datasets and seven fidelity and utility metrics to thoroughly assess these methods. Notably, we find that the common practice of selecting downstream classifiers based on the highest accuracy on the sensitive test set not only violates DP but also overestimates the utility scores. DPImageBench corrects for these mistakes. (3) Platform. Despite the wide variety of methods and evaluation protocols, DPImageBench provides a standardized interface that accommodates current and future implementations within a unified framework. With DPImageBench, we have several noteworthy findings. For example, contrary to the common wisdom that pretraining on public image datasets is usually beneficial, we find that the distributional similarity between pretraining and sensitive images significantly impacts the performance of the synthetic images and does not always yield improvements. In addition, adding noise to low-dimensional features, such as the high-level characteristics of sensitive images, is less affected by the privacy budget compared to adding noise to high-dimensional features, like weight gradients. The source code is available.111https://github.com/2019ChenGong/DPImageBench
1. Introduction
Privacy-preserving synthetic data aims to generate artificial data that maintains the properties of real data, enabling data sharing within and across organizations while minimizing privacy risks (Hu et al., 2024; Lin, 2022). Differentially private (DP) dataset synthesis (Hu et al., 2024; Du and Li, 2024) provides a theoretical guarantee to quantify privacy leakage of real data using the synthetic dataset. In recent years, there has been significant development of DP dataset synthesis methods across various domains, including images (Li et al., 2024; Lin et al., 2024, 2025; Dockhorn et al., 2023), tabular data (Cai et al., 2021; McKenna et al., 2022), text (Yue et al., 2023; Xie et al., 2024; Yu et al., 2023), networking data (Lin et al., 2020; Yin et al., 2022; Sun et al., 2024) and more.
Among them, DP image synthesis has witnessed impressive progress (Lin et al., 2024, 2025; Li et al., 2024; Jiang et al., 2023; Tsai et al., 2024) but lacks a comprehensive understanding of the strengths and weaknesses of existing synthesizers. While there are several studies summarizing advances in this field, they either focus on discussions (Chen et al., 2023; Hu et al., 2024) and do not perform much evaluation, or only cover subfields (Wei et al., 2023; Ganev et al., 2024). This paper focuses on benchmarking DP image dataset synthesis, addressing several challenges:
@@ 1. Introduction
Evaluation: The lack of unified metrics makes it difficult to evaluate the quality of synthetic images. Existing methods, such as PE (Lin et al., 2024) and PrivImage (Li et al., 2024), both utilize FID (Heusel et al., 2017) for fidelity evaluation. However, FID does not assess whether the generated images align with human perception, as described in Section 5.3.2. Even worse, the standard practice that evaluates synthetic image quality by tuning classifiers on the sensitive test set to select optimal parameters (used in several methods (Tsai et al., 2024; Chen et al., 2020)) is incorrect: they both violate DP and inflate the results.
@@
Contributions. Without a clear understanding of existing methods and an easy-to-use benchmarking platform, it is challenging for the community to make meaningful and reproducible advancements in this field. To bridge this gap, we establish DPImageBench, a benchmark focused on DP image synthesis. DPImageBench addresses the above challenges by making the following contributions:
@@
Taxonomy. We taxonomize methods from two orthogonal dimensions, covering twelve state-of-the-art DP image synthesis algorithms. (1) Private Data Perspective. Based on where noise is added to satisfy DP, we categorize methods into (a) Input-level: adding noise to (some format of) the sensitive images; (b) Model-level: incorporating noise into the gradient or loss during model training; and (c) Output-level: introducing noise during the synthesis process. (2) Public Data Perspective. Existing works (Li et al., 2024; Liu et al., 2023) use publicly available data from open platforms without privacy concerns to enhance DP image synthesis. We categorize them into three types based on how the public data is utilized: leveraging pretrained models, public datasets without selection, and public datasets with selection. Besides, we argue that the state-of-the-art method DP-Promise (Wang et al., 2024) does not strictly satisfy DP in Appendix E to prevent others from wrongly using it.
@@
Platform. We establish a platform, DPImageBench, with the following contributions: (1) Fair comparison. DPImageBench standardizes the public images, hyper-parameters, and model architectures to ensure fair comparisons. DPImageBench offers an interface to pretrain the synthesizer, even for the ones that were not originally designed to use a public dataset for pretraining. (2) Straightforward implementation. DPImageBench features an abstract interface compatible with any synthesis algorithm. With its modular design, DPImageBench provides a flexible tool that enables users to select public and sensitive datasets, and to integrate new algorithms by adding functional code to the relevant modules. Since DPImageBench isolates the parameter configurations from implementations, users typically only need to modify the configuration files to tune algorithms. (3) Open Source. DPImageBench is open-sourced. It serves as a starting point for the implementation and evaluation of future methods as well.
@@
Evaluation. We have conducted comprehensive evaluations, and some key highlights include: (1) Datasets. DPImageBench implements synthesis methods on seven sensitive datasets and two public dataset. (2) Fidelity. We establish metrics to assess synthetic images based on similarity, diversity, novelty, and alignment with human perception (Xu et al., 2024). (3) Utility. We correct the common evaluation mistakes of using sensitive test images for hyper-parameter tuning. Instead, DPImageBench adopts the standard of using distinct training, validation, and test sets for utility evaluations. Downstream classifier training is performed on the training set, hyper-parameter selection is done on the validation set, and final results are reported on the test set. DPImageBench provides two choices of validation set: (1) the synthetic images, and (2) images having the same distribution as the test set but not overlapping with it. In the latter, DPImageBench adds noise to the validation results to satisfy DP, which is detailed in Section 5.3.4.
@@
Insights to new techniques. The proposed taxonomies and straightforward implementations enable us to analyze the potential of combining different methods and bring insight into developing new technology. We consider some methods mostly in isolation, though they can be effectively combined. For example, the public dataset selection method from PrivImage (Li et al., 2024) and creating shortcuts for DP training in DP-FETA (Li et al., 2025), as described in Section 4, can be paired with other methods. We present details in Section 7.4.
@@
Pretraining with public datasets does not always benefit DP image synthesis. The effectiveness of pretraining depends heavily on the choice of public dataset and the pretraining strategy (i.e., conditional or unconditional pretraining). Conditional pretraining yields better synthetic performance than unconditional pretraining. A pretraining dataset similar to the sensitive images can enhance synthetic performance.
The performance of adding noise to low-dimensional features, such as high-level characteristics of sensitive images in DP-MERF (Harder et al., 2021) and training loss in DP-Kernel (Jiang et al., 2023), is less sensitive to privacy budgets compared to adding noise to high-dimensional features, like weight gradients using DP-SGD. The former methods perform better than the latter under a low privacy budget.
The combinations of algorithmic improvements, such as integrating DP-FETA (Li et al., 2025) into other methods, do not consistently enhance synthetic performance. This inconsistency underscores the need for further research to effectively leverage the strengths of diverse algorithms.
2. Preliminaries
2.1. Privacy Concepts and Mechanisms
DPImageBench concentrates on image synthesis under differential privacy (DP) constraint. DP (Dwork et al., 2006) provides mathematical guarantees that quantify and limit the amount of information one can derive from the algorithm’s output.
Definition 0 (()-DP (Dwork et al., 2006)).
@@ 2.1. Privacy Concepts and Mechanisms
The privacy budget quantifies how much information can be exposed by the algorithm . A smaller value of implies that the algorithm provides stronger privacy guarantees. The can be intuitively understood as failure probability (Dwork et al., 2006). Given two datasets and , they are considered neighbors (denoted as ) if we can obtain one from the other by adding, removing, or replacing a single image. We define the sensitivity (intuitively, a stability measure) of a function as . As presented in Table 12 of Appendix A, two widely used definitions for neighboring datasets exist in existing DP image synthesis: the ‘replace-one’ (bounded DP) and ‘add-or-remove-one’ (unbounded DP) definitions. Specifically, given two datasets and , in the ‘replace-one’ definition, neighboring datasets are defined by replacing one image sample, i.e., for image samples and . In the ‘add-or-remove-one’ definition, neighboring datasets are constructed by adding or removing a single image sample, e.g., . We show in Appendix A that all methods that used the replace-one definition can be adapted to add-or-remove-one without affecting their privacy guarantees. As a result, all methods can be compared fairly under the same privacy notion.
@@
is a multi-dimensional random variable drawn from an i.i.d. normal distribution with a mean of 0 and a variance of , where is a hyper-parameter known as the noise multiplier (Dwork et al., 2006).
DP-SGD. To train machine learning models with DP guarantees, DP-SGD (Abadi et al., 2016) extends standard SGD by controlling the gradient contributions from individual data and applying the Gaussian mechanism. Specifically, in each training iteration, DP-SGD samples a mini-batch from the whole training data and bounds the sensitivity of each gradient using a clipping operation: where indicates the original gradient and is a hyper-parameter. The ‘’ operation scales the gradient norm to be at most , and the model parameters are then updated with:
where is the learning rate, denotes the gradient computed from the loss function for the sample , and refers to a mini-batch sampled from the dataset using Poisson sampling with a rate . Besides, is i.i.d. Gaussian noise with variance .
As mentioned above, DP-SGD uses only one batch of the sensitive dataset for each iteration of the model parameter update. However, the training process involves many iterations of DP-SGD to ensure that the target model converges. Composition theorems (privacy accounting), e.g., Rényi DP (Mironov et al., 2019), can be used to account for the privacy cost for each iteration and derive the final values of (Mironov et al., 2019).
2.2. Image Generative Models
We introduce three image synthesizers: (1) diffusion models (DMs), including standard DMs (Ho et al., 2020) and latent DMs (Rombach et al., 2022), (2) Generative Adversarial Networks (GANs) (Arjovsky et al., 2017), and (3) foundational APIs (Lin et al., 2024).
Diffusion Models. DMs (Ho et al., 2020) consist of two processes: (1) The forward diffusion process that progressively adds noise to a clean image until it becomes close enough to pure random noise, generating a sequence of increasingly noisy images , where is the number of noising steps. (2) The reverse process that progressively denoises a random noise to a clean image via a trainable neural network. In the forward diffusion process, the transition between adjacent noisy images, , follows a multi-dimensional Gaussian distribution,
@@ 2.2. Image Generative Models
The final objective of DMs is to train a network to predict the added noise at each step, as defined in (Ho et al., 2020).
@@
Latent Diffusion Models (LDMs). LDMs (Rombach et al., 2022) are a variant of DMs that operate in a latent space instead of directly in pixel space. LDMs first use a pretrained autoencoder to map images into a lower-dimensional latent representation. The same diffusion and denoising steps discussed before proceed in this compressed latent space rather than on raw pixels. After performing the reverse diffusion steps in latent spaces, the decoder reconstructs final high-resolution images. LDMs operate in a lower-dimensional space, enabling more efficient training and sampling than standard pixel-based DM.
Stable diffusion is a representative example of LDMs that generate images from textual descriptions or class labels. DP-LDM (Liu et al., 2023) and DP-LoRA (Tsai et al., 2024) leverage stable diffusion (Rombach et al., 2022) as the synthesizer.
Generative Adversarial Nets (GAN). GAN consists of two networks: a generator Gen and a discriminator Dis (Arjovsky et al., 2017; Goodfellow et al., 2014). The Gen receives a random noise vector and outputs an image. The Dis receives an image and outputs a score, which indicates how real the input image is. The Gen is trained to deceive the discriminator, and the Dis is trained to distinguish whether its input image comes from the true dataset or the Gen. Both networks are trained simultaneously, constantly improving in response to each other’s updates. Overall, the objective function of GAN is defined as,
@@
Foundational APIs. Foundation APIs are either cloud-based services (e.g., Dall-E (Ramesh et al., 2022)) or local software libraries (e.g., Stable Diffusion (Rombach et al., 2022)) that provide simple API access to pretrained models capable of generating high-quality synthetic images. The underlying model can be any generative model, including DMs and GANs.
2.3. DP Image Synthesis
Various sensitive image datasets, such as medical images (Bándi et al., 2019) and face images (Liu et al., 2015), face the risk of privacy leakage. Attackers may infer information about the training dataset from the synthetic images (Matsumoto et al., 2023). DP image synthesis aims to generate artificial images that closely resemble real data while ensuring the privacy protection of the original dataset. DP quantifies leakage to infer the training dataset using the synthetic data. They enable organizations to share and use synthetic images, reducing privacy concerns.
3. Overview of DPImageBench
We introduce DPImageBench, a modular toolkit designed to evaluate DP image synthesis algorithms in terms of utility and fidelity. We envisage DPImageBench for the following four purposes.